OpenStudy (anonymous):

1. Complete the following ordered pairs for the equation. x + 4y = 10 (-2, _) (_, 2) (6, _)

OpenStudy (ohmaigosshh):

(-2, ) x + 4y = 10. -2 + 4y = 10. Add 2 to both sides. 4y = 12 Divide both sides by 4 y = 3 .... so for the first one it is (-2,3)

OpenStudy (ohmaigosshh):

Second one... ( , 2) x + 4y = 10 x + 4(2) = 10 x + 8 = 10 Subtract 8 from both sides x = 2 ... so it will be (2,2)

OpenStudy (ohmaigosshh):

Third one... (6, 0) x + 4y = 10 6 + 4y = 10 subtract 6 from both sides 4y = 4 Divide both sides by 4 y = 1.... so it will be (6, 1)

OpenStudy (anonymous):

thanks for show me how to do it i think i got how to do this now

OpenStudy (ohmaigosshh):

Your Welcome :)